# mycap[Listing][List of Listings]

### v1

@@ 4.1 Experimental Setup
We evaluated a range of prominent and representative models on ACVUBench, encompassing audio-visual, visual-only, and audio-only architectures, to comprehensively assess performance across different modalities. Audio-Visual MLLMs: We evaluated Gemini 1.5 ProTeam et al. (2024), Video-SALMONN(13B)Sun et al. (2024b), VideoLLaMA2(7B)Cheng et al. (2024), and PandaGPT(13B)Su et al. (2023). These models received both video and audio inputs. Visual MLLMs: We evaluated GPT-4oHurst et al. (2024), Qwen-2-VL(7B)Wang et al. (2024), LLaVA-video(7B)Zhang et al. (2024), InternVL2(76B and 8B)OpenGVLab (2024), VILA-1.5(8B)Chen et al. (2024), and VideoLLaVA(7B)Lin et al. (2023). These models received video input with the audio track removed. Audio MLLM: We evaluated SALMONN(13B)Tang et al. (2024) using only the audio track extracted from the videos.
@@ 4.2 Results
Audio is crucial in ACVUVBench.
To support this, we choose the most competitive audio-visual model, Gemini 1.5 Pro, to perform ablation studies as shown in Table 5. This demonstrated that our benchmark relies heavily on audio modality, as its performance drops significantly when missing any of the two modalities.
@@
Table 5: Gemini-1.5-pro(2024)’s performance when provided different modalities.
Modalities Audio-Visual Visual Only Audio Only Accuracy     78.34     62.78     54.10
To further study this issue, we introduce the confidence thresholded accuracy in section 4.3 to indicate how much the model is confused, hence inferring how much the model is suffering from hallucination.
@@
Models GPT rank Rouge-L MCQ Acc Audio Visual MLLMs Gemini 1.5 Pro (2024)     3.331     0.108     78.34 VideoLLaMA2 (7B) (2024)     1.839     0.062     44.90 video-SALMONN (13B) (2024b)     2.032     0.080     38.33 PandaGPT (13B) (2023)     1.384     0.035     25.38 Visual MLLMs Qwen2-VL (7B) (2024)     2.233     0.074     58.38 LLaVA-Video (7B) (2024)     2.217     0.099     56.52 InternVL2 (76B) (2024)     2.022     0.053     52.62 InternVL2 (8B) (2024)     1.959     0.052     45.90 VILA-1.5 (8B) (2024)     2.079     0.074     44.48 VideoLLaVA (7B) (2023)     1.024     0.022     33.14 Audio MLLM SALMONN (13B) (2024)     1.758     0.053     36.48
4.3 Confidence-Thresholded Accuracy
Standard accuracy provides insufficient insights for MCQ, particularly in ACVUBench since two different non-text modalities are involved. As a result, audio-visual models are likely to have more hallucinations Sun et al. (2024a). A model suffering from hallucination is more likely to produce uncertain predictions, and answer questions by random guessing. Therefore, to distinguish informed decision-making from guesses in accuracy metrics, a dedicated confidence-thresholded accuracy (CTA) metric is necessary.
We use the first token probability to compute CTA by instructing the model to output a single letter corresponding to its chosen option. We take the logits corresponding to the set of possible choices (e.g. A, B, C and D) from the output distribution of the first generated token and normalise them to a valid distribution, as shown in Eqn. (1).
$P(c)=\frac{{\exp}(l_{c})}{\sum_{c^{\prime}\in\mathcal{C}}{\exp}(l_{c^{\prime}})},$ (1)
where $c$ is the token of the choice, $l_{c}$ is the logit value of the choice token $c$, and $\mathcal{C}$ is the set of all possible choice tokens. Then we compute the difference between the probability assigned to the correct answer and the probability assigned to the most likely incorrect answer (i.e., the maximum probability among the incorrect options):
$\Delta P=P(c_{\text{ref}})-\max_{c\in\mathcal{C}\setminus\{c_{\text{ref}}\}}P(c),$ (2)
where $c_{\text{ref}}$ is the correct answer token. Finally, we introduce the threshold, $p\in(-1,1)$, where a prediction is considered correct only if the probability difference exceeds the given threshold:
$\text{CTA}=\frac{1}{N}\sum_{n}^{N}\mathbbm{1}(\Delta P_{n}\geq p),$ (3)
where $\mathbbm{1}(\cdot)$ is the indicator function and $N$ is the total number of test samples.
Figure 3: CTA curve for SALMONN, video-SALMONN and Qwen2-VL-7B measured using all MCQs in ACVUBench.
We plot the variation of CTA against threshold $p$ for three models in Figure 3. A steeper curve means that the model is performing more random guesses. As shown, Qwen-2-VL-7B relies significantly less on random guessing compared to others. We believe this is due to Qwen2 being a more recently developed LLM backbone than Vicuna-v1.5 in SALMONN and video-SALMONN, hence suffering much less from hallucination. In addition, commonly used confidence measures, such as the expected calibration error (ECE), are unable to indicate the level of random guesses. As a result, the three models yielded low ECE values between 0.049 and 0.122, necessitating the use of CTA.
@@ 5 Conclusion
In this paper, we present ACVUBench, a comprehensive benchmark designed to advance audio-centric video understanding. ACVUBench provides a diverse suite of tasks focused on audio understanding and audio-visual alignment, supported by a meticulously curated dataset including both human-annotated and Gemini-augmented subsets. An extensive evaluation of models highlights the importance of audio information in complete video comprehension. We believe ACVUBench is valuable for the research community, fostering innovation and progress in audio-centric video understanding and paving the way for more robust and comprehensive multimodal learning systems.
@@
While we endeavor to ensure a wide range of annotators, it is inevitable that certain annotator bias still exists in data annotation, such as preference towards visual information and bias due to the first language. Moreover, same as many other MCQ-based benchmarks, the MCQ parts of ACVUBench only explored 4 option cases, leaving other numbers of choices unexplored.
@@ Appendix A Example Videos and Tasks
This section provides illustrative examples of the six multiple-choice question tasks in ACVUBench, the video data and corresponding task is in Figure 4 . Each figure showcases a video frame and the corresponding annotated question with answer options.
@@
Figure 4: Six MCQ format task examples in ACVUBench.
@@ Appendix E Annotation Guidelines and Examples
This section details the annotation process for creating question-answer pairs for the AV-Human. Annotators were provided with 700 videos and tasked with creating three question-answer pairs per video, totaling 2,100 annotations. The annotation process followed these steps:
@@ Appendix F Gemini-Generated Annotation Review Process
This section details the human review process for the 11,742 Gemini-generated question-answer pairs. Annotators were tasked with ensuring the generated content met specific quality criteria, focusing on question relevance, option validity, answer accuracy, and adherence to video content without reliance on external knowledge or common sense reasoning. The review process for each question-answer pair involved the following steps:
@@
Our human review stage was critical for mitigating biases and ensuring annotation quality. Approximately 40% of the 11,742 generated QA pairs required modification, ranging from minor rephrasing to complete rewriting. Common categories of issues encountered during the review process are detailed below, along with illustrative examples:
@@ Appendix G Data Availability and Copyright
Our benchmark ACVUBench is constructed using publicly available videos from YouTube. We adhere to YouTube’s copyright policies and terms of service. ACVUBench contains only links to these public videos; we do not host or distribute copies of the video content itself. This ensures compliance with copyright regulations and allows for easy access to the data while respecting intellectual property rights. Users of ACVUBench are responsible for complying with YouTube’s terms of service when accessing the linked videos.
@@
Gemini 1.5 Pro: Gemini 1.5 Pro was employed solely for data augmentation purposes, specifically to generate a larger set of question-answer pairs for our benchmark. In accordance with Google’s terms of service, we acknowledge that Google retains all rights to generate similar content for other users. We do not claim ownership over the Gemini-generated content. We also acknowledge our responsibility for the use of this generated content, both within our research and by anyone with whom we share the benchmark data. While we have used discretion in relying on the generated content, users of ACVUBench should be aware of its automated origin. As required by the API Terms, we will comply with applicable laws regarding the use of generated content, including providing attribution when necessary.
@@ Appendix I Model Prompting Details
Prompt 1 is used for evaluating models on the multiple-choice question tasks in ACVUBench. This prompt was provided to each model.

### v3

@@ 5.1 Experimental Setup
We evaluated a range of prominent and representative models on AVUT, encompassing audio-visual, visual-only, and audio-only architectures, to comprehensively assess performance across different modalities. Audio-Visual MLLMs: We evaluated Gemini 1.5 Pro Team et al. (2024), Video-SALMONN(13B) Sun et al. (2024), VideoLLaMA2(7B) Cheng et al. (2024), and PandaGPT(13B) Su et al. (2023). These models received both video and audio inputs. Visual MLLMs: We evaluated GPT-4o Hurst et al. (2024), Qwen-2-VL(7B) Wang et al. (2024), LLaVA-video(7B) Zhang et al. (2024), InternVL2(76B and 8B) OpenGVLab (2024), VILA-1.5(8B) Chen et al. (2024), and VideoLLaVA(7B) Lin et al. (2023). These models received video input with the audio track removed. Audio MLLM: We evaluated SALMONN(13B) Tang et al. (2024) using only the audio track extracted from the videos.
@@ 5.2 Results
AVUT is designed to be a multi-modal benchmark where audio plays a crucial role.
To support this, we choose the most competitive audio-visual model, Gemini 1.5 Pro, to perform ablation studies as shown in Table 5. We compared the model’s performance across several input configurations: full audio-visual, video-only, audio-only, and a detailed transcription setting. The "detailed transcription" includes speaker identification, corresponding ASR transcripts, and start/end timestamps for each utterance. This result demonstrated that our benchmark relies heavily on audio modality, as its performance drops significantly when missing any of the two modalities. While providing detailed transcriptions with speaker diarization and timestamps improves performance compared to video-only, it still falls short of the performance achieved with full audio. This suggests that beyond the transcribed words themselves, the rich information encoded in audio, such as intonation, prosody, and overlapping speech, is crucial for understanding the video content and is not fully captured even by detailed textual transcriptions. Transcription annotation examples are shown in Appendix K.
@@
Modalities Audio-Visual Transcription Visual-Only Audio-Only Accuracy (%)     78.34     69.23     62.78     54.10
Table 5: Gemini 1.5 Pro’s performance when provided different modalities.
@@
Models GPT rank Rouge-L MCQ Acc Audio Visual MLLMs Gemini 1.5 Pro (2024)     3.331     0.108     78.34 VideoLLaMA2 (7B) (2024)     1.839     0.062     44.90 video-SALMONN (13B) (2024)     2.032     0.080     38.33 PandaGPT (13B) (2023)     1.384     0.035     25.38 Visual MLLMs Qwen2-VL (7B) (2024)     2.233     0.074     58.38 LLaVA-Video (7B) (2024)     2.217     0.099     56.52 InternVL2 (76B) (2024)     2.022     0.053     52.62 InternVL2 (8B) (2024)     1.959     0.052     45.90 VILA-1.5 (8B) (2024)     2.079     0.074     44.48 VideoLLaVA (7B) (2023)     1.024     0.022     33.14 Audio MLLM SALMONN (13B) (2024)     1.758     0.053     36.48
@@
5.3 Evaluating Text Shortcut on Benchmarks
Benchmarks Text-Only Accuracy (%) $\downarrow$ Cycled Accuracy (%) $\downarrow$     GPT-4o     Gemini 1.5 Pro     GPT-4o     Gemini 1.5 Pro VideoVista (2024c)     68.00     62.63     41.25     38.75 TemporalBench (2024)     65.42     59.37     73.57     52.98 NExT-QA (2021)     47.57     48.95     23.22     26.37 MLVU (2024)     42.96     40.98     19.38     12.81 MotionBench (2025)     41.24     36.37     16.17     11.86 Video-MME (2024a)     41.22     45.22     16.03     13.05 MVBench (2024a)     32.24     39.03     10.84     5.70 Perception Test (2023)     29.20     32.67     4.68     6.02 AVUT     25.32     29.61     1.89     4.50
Table 8: Benchmark Text Shortcut Result. "Text-Only Accuracy" denotes accuracy with text input only. "Cycled Accuracy" denotes accuracy requiring consistent correct answers across four option permutations (ABCD, BCDA, CDAB, DABC). Text shortcut is tested with both GPT-4o and Gemini 1.5 Pro to avoid model bias.
To demonstrate the prevalence of text shortcut and the effectiveness of our mitigation strategy, we evaluated several video understanding benchmarks alongside AVUT. We used GPT-4o as our test model, assessing its performance in two conditions: (1) using the Text-Only Accuracy (only the question text is given) and (2) using the Cycled Accuracy metric (requiring correct answers across all four permutations).
As shown in Table 8, existing benchmarks suffer from text shortcut, with both GPT-4o and Gemini-1.5-pro achieving surprisingly high accuracy without accessing any video. In particular, VideoVista had a text-only accuracy of 68.00% with GPT-4o, and the audio-visual benchmark, Video-MME had 41.22%, which are all far above the 25% accuracy of random guessing. This highlights the ubiquity of the problem and the potential for misleading performance evaluations. In contrast, AVUT approaching the expected random guessing accuracy on four-option multiple-choice questions with text-only input on both GPT-4o and Gemini-1.5-pro. Furthermore, AVUT also achieves a cycled accuracy of 1.89% with GPT-4o that is significantly lower than others by large margins, confirming the effectiveness of our filtering mechanism in mitigating text shortcut. As a result, AVUT provides a more reliable and rigorous assessment of the actual capabilities of video understanding.
@@ 6 Conclusion
In this paper, we present AVUT, a text-shortcut-free benchmark designed to advance audio-centric video understanding. AVUT provides a diverse suite of tasks focused on audio understanding and audio-visual alignment, supported by a meticulously curated dataset including both human-annotated and Gemini-augmented subsets. An extensive evaluation of models highlights the importance of audio information in complete video comprehension. We believe AVUT is valuable for the research community, fostering innovation and progress in audio-centric video understanding and paving the way for more robust and comprehensive multimodal learning systems.
@@
While we endeavor to ensure a wide range of annotators, it is inevitable that certain annotator bias still exists in data annotation, such as preference towards visual information and bias due to the first language. Moreover, same as many other MCQ-based benchmarks, the MCQ parts of AVUT only explored 4 option cases, leaving other numbers of choices unexplored.
@@ Appendix A Example Videos and Tasks
This section provides illustrative examples of the six multiple-choice question tasks in AVUT, the video data and corresponding task is in Figure 3 . Each figure showcases a video frame and the corresponding annotated question with answer options. And 18 audio-centric video domains are shown in Figure 2.
Figure 2: The 18 video domains of AVUT from which the videos are collected.
@@
Figure 3: Six MCQ format task examples in AVUT.
@@ Appendix E Annotation Guidelines and Examples
This section details the annotation process for creating question-answer pairs for the AV-Human. Annotators were provided with 700 videos and tasked with creating three question-answer pairs per video, totaling 1,734 annotations. The annotation process followed these steps:
@@ Appendix F Gemini-Generated Annotation Review Process
This section details the human review process for the 9,875 Gemini-generated question-answer pairs. Annotators were tasked with ensuring the generated content met specific quality criteria, focusing on question relevance, option validity, answer accuracy, and adherence to video content without reliance on external knowledge or common sense reasoning. The review process for each question-answer pair involved the following steps:
@@
Our human review stage was critical for mitigating biases and ensuring annotation quality. Approximately 40% of the 9,875 generated QA pairs required modification, ranging from minor rephrasing to complete rewriting. Common categories of issues encountered during the review process are detailed below, along with illustrative examples:
@@ Appendix G Data Availability and Copyright
Our benchmark AVUT is constructed using publicly available videos from YouTube. We adhere to YouTube’s copyright policies and terms of service. AVUT contains only links to these public videos; we do not host or distribute copies of the video content itself. This ensures compliance with copyright regulations and allows for easy access to the data while respecting intellectual property rights. Users of AVUT are responsible for complying with YouTube’s terms of service when accessing the linked videos.
@@
Gemini 1.5 Pro: Gemini 1.5 Pro was employed solely for data augmentation purposes, specifically to generate a larger set of question-answer pairs for our benchmark. In accordance with Google’s terms of service, we acknowledge that Google retains all rights to generate similar content for other users. We do not claim ownership over the Gemini-generated content. We also acknowledge our responsibility for the use of this generated content, both within our research and by anyone with whom we share the benchmark data. While we have used discretion in relying on the generated content, users of AVUT should be aware of its automated origin. As required by the API Terms, we will comply with applicable laws regarding the use of generated content, including providing attribution when necessary.
@@ Appendix I Model Prompting Details
Prompt 1 is used for evaluating models on the multiple-choice question tasks in AVUT. This prompt was provided to each model.
@@
Appendix K Transcription Generation with Gemini
This section provides examples of the detailed transcriptions generated using Gemini 1.5 Pro for AVUT. These transcriptions include speaker identification, corresponding ASR transcripts, and start/end timestamps for each utterance. The inclusion of detailed transcriptions allows us to investigate the extent to which textual representations of audio can substitute for the original audio in video understanding tasks.
"start_time": 0.0,
"end_time": 2.0,
"speaker": "Speaker 1",
"characteristic": "female with long brown hair, wearing a blue dress with flowers",
"transcript": "I heard that you brought your parents to the Oscars was that fun?"
"start_time": 2.0,
"end_time": 6.0,
"speaker": "Speaker 2",
"characteristic": "female with blonde hair, wearing a blue jacket with flowers",
"transcript": "I did I did my dad actually called me and sort of invited himself."
Listing 1: Example Transcription